OpenStudy (anonymous):

Teesha is in the French club. There are 31 students in the club. The French teacher will pick two students at random to guide visiting students from France. What is the probability that Teesha will be picked as a guide? Hint: Teesha must be picked first or second. A. B. C. D.

OpenStudy (anonymous):

\[Teesha is \in the French club. There are 31 students \in the club. The French teacher will pick two students at random \to guide visiting students from France. What is the probability that Teesha will be picked as a guide? Hint: Teesha must be picked first or second. A. \frac{ 1 }{3 } B. C. D.\]

OpenStudy (kropot72):

This can be solved by using the hypergeometric distribution as follows: \[P(Teesha)=\frac{\left(\begin{matrix}30 \\ 1\end{matrix}\right)}{\left(\begin{matrix}31 \\ 2\end{matrix}\right)}=\frac{30\times 2}{31\times 30}\]